OpenStudy (anonymous):

please help me go through the steps in solving this 1. Rewrite 2x + 3y = 9 in slope-intercept form and then graph the line.

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

isolate y

OpenStudy (anonymous):

\[2x + 3y = 9\\ 3y = 9 -2x\\ y=\frac{9-2x}{3}\]

OpenStudy (anonymous):

thanks